OpenStudy (needshelpdfdf):

Solve for x. x2 − 2x = 0

OpenStudy (eliesaab):

Same \[ x^2-2x=x(x-2) \]

OpenStudy (needshelpdfdf):

I don't know

OpenStudy (will.h):

Find the GCF the GCF here is x so x(x - 2) = 0 So either x = 0 or x - 2 = 0 x = 2

OpenStudy (eliesaab):

x=0 and x=2

OpenStudy (needshelpdfdf):

oh so it 0,2

OpenStudy (eliesaab):

Yes

OpenStudy (needshelpdfdf):

+ right

OpenStudy (eliesaab):

yes